OpenStudy (anonymous):

Please Help : 4(x + 2) = 6^x-7

OpenStudy (anonymous):

is it 4(x + 2) = (6^x)-7 , or , 4(x + 2) = 6^(x-7)

OpenStudy (anonymous):

6^x = 4 x+15

OpenStudy (anonymous):

by multiplying amd simplifying

OpenStudy (anonymous):

I mean no!!1

OpenStudy (anonymous):